OpenStudy (anonymous):

y=-x+2 and y=-2x+6 ordered pairs ?

OpenStudy (anonymous):

Same thing again, you can use substitution, elimination, or matrices. For this, just use substitution. -x + 2 = -2x + 6 x = 4 Then plug it back in for y. y = -x + 2 y = -4 + 2 y = -2 (4, -2)